Grade upto college level Physical Chemistry

A gas mixture of 3.67 litres of ethylene and methane on complete combustion at 25oC produces 6.11 litres of CO2. Find out the amount of heat evolved on burning one litre of the gas mixture. The heats of combustion of ethylene and methane are -1423 and -891 kJ mol-1 at 25oC.

Shane Macguire
12 Years agoGrade upto college level

2 Answers

Deepak Patra
12 Years ago
Hello Student,
Please find the answer to your question
Combustion of C2H4 and Ch4 takes place as follows :
C2H4 + 3O2 → 2CO2 + 2H2O
1 vol 2 vol.
CH4 + 2O2 → CO2 + 2H2O
1 vol. 1 vol
Let the vol. of CH4 in mixture = x l
∴ Vol. of C2H4 in the mixture = (3.67 - x) l
Vol. of CO2 produced by x l of CH4 = x l and
Vol. of CO2 produced by (3.67 - x) l of C2H4 = 2(3.67 - x) l
∴ Total vol. of CO2 produced = x + 2 (3.67 - x)
Or 6.11 = x + 2(3.67 - x) or x = 1.23 l
∴ Vol. of CH4 in the mixture = 1.23 l
and Vol. of C2H4 in the mixture = 3.67 – 1.23 = 2.44 l
Vol. of CH4 per litre of the mixture = 1.23/3.67 = 0.335 l
Vol. of C2H4 per litre of the mixture = 2.44/3.67 = 0.665 l
Now we know that volume of 1 mol. Of any gas at 25°C (298 K) = 22.4 * 298/273 = 24.45 l
[∵ Volume at NTP = 22.4L]
Heat evolved due to combustion of 0.335 l of CH4 = - 0.335 * 891/24.45 = - 12.20 kJ [given, heat evolved by combustion of 1 l = 891 kJ]
Similarly, heat evolved due to combustion due to combustion of 0.665 l of C2H4
= - 0.665 * 1423/24.45 = - 38.70 kJ
∴ Total heat evolved = 12.20 + 338.70 = 50.90 kJ

Thanks
jitender
askIITians Faculty
nkc
8 Years ago
make the eqn of combustion and then use addition aur subtraction to get desired eqn and then apply same operation on heat given